OpenStudy (anonymous):

Margaret is planning a rectangular garden. Its length is 4 feet less than twice its width. Its area is 170ft^2. what are the dimensions of the garden?

OpenStudy (mathstudent55):

The area of a rectangle is A = LW, where A = area, L = length, W = width The width is W The length is 4 feet less than twice the width twice the width = 2W four feet less than twice the width = 2W - 4 L = 2W - 4 A = LW, but A = 170, so 170 = LW L = 2W - 4 Since the second equation is already solved for L, just plug that into the first equation and solve for W. Then plug in W in the second equation and solve for L

OpenStudy (anonymous):

170=(2w-4)w once you get here do you distribute?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

yes, distribute

OpenStudy (anonymous):

170=(2w-4)w 170=2w^2-4w 0=2w^2-4w-170 I dont think im doing this right

OpenStudy (mathstudent55):

It looks good to me.

OpenStudy (mathstudent55):

Now try to factor the right side.

OpenStudy (anonymous):

can you help me with that, im not sure how

OpenStudy (mathstudent55):

First, divide both sides by 2

OpenStudy (mathstudent55):

also, switch sides

OpenStudy (mathstudent55):

2w^2 - 4w - 170 = 0 Divide both sides by 2

OpenStudy (anonymous):

okay so 2w^2-4w-170=0 w^2-2w-85=0?

OpenStudy (mathstudent55):

Right. Now this kind of factoring involves simply finding two numbers that multiply to -85 and add to -2

OpenStudy (mathstudent55):

-85 = -85 x 1 -85 + 1 = -84 -85 = 85 x -1 85 - 1 = 84 -85 = -17 x 5 -17 + 5 = -12 -85 = 17 x (-5) 17 - 5 = 12

OpenStudy (mathstudent55):

As you can see, there are no two such numbers. This cannot be factored. We now use the quadratic equation. Let's move to a drawing.

OpenStudy (mathstudent55):